# Splat Feature Solver

### v1

@@ Introduction
Here, we start with the original loss function $\mathcal{L}(x)$, and calculate a surrogate loss function $\mathcal{J}(x)$. The norm $||\cdot||$ represents any convex loss function such as L1, L2, or a Huber loss. From Property.2, we apply the Jensen’s inequality. This means the surrogate loss function $\mathcal{J}$ is larger than $\mathcal{L}$ and serves as an upper bound for the true loss. For the sake of argument, we consider a special case where $||\cdot||$ is the L2 loss. If $\mathcal{J}$ attains a minimum, it occurs where the gradient is zero, as shown in Equ.12. Therefore, we obtain an optimal solution on the surrogate loss function $\mathcal{J}$.
$\frac{\partial\mathcal{J}}{\partial x_{j}}$ $=\sum_{i}A_{ij}\left(x_{j}-B_{i}\right)$ (11)
$\frac{\partial\mathcal{J}}{\partial x_{j}}=0$ $\quad\Rightarrow\quad x^{\prime}_{j}=\frac{\sum_{i}A_{ij}B_{i}}{\sum_{i}A_{ij}}$ (12)
We now define $\beta$, which intuitively measures the dispersion of the lifted feature along a viewing ray at the global optimal lift. Suppose the optimal lift yields a solution $\hat{x}$. We define $\beta$ as shown in the following Equ.14.
@@
Notice that usually $x^{\prime}\neq\hat{x}$. Therefore, $x^{\prime}$ is not optimal in $\mathcal{J}$. By Equ.10, we have the following equation under the L2 loss:
[…]
@@
From the analysis, we can conclude that the methods in(Dou et al. 2024; Cheng et al. 2024) are also bounded by the $1+\beta$ term. The method (Chacko et al. 2025) is bounded when a $l_{\infty}$ norm function is applied. The approach in (Joseph, Amrutur, and Bhatnagar 2024) is bounded only when the input observations are unit vectors. (Jun-Seong et al. 2025) does not provide a theoretical bound, as it selects the top-K splats based solely on depth ordering.
Regularizer
@@
Tikhonov Guidance
According to property.4 and inspired by (A. and V. 1977; A. 1963), we observe that emphasizing diagonal dominance helps mitigate noise from linear system A. Therefore, we propose a carefully designed regularizer applied during solving without sacrificing the RGB rendered result. Specifically, we first convert Equ.7 into a fully diagonal row-sum version. . Briefly speaking, stabilizing the linear system $A^{T}A$ involves avoiding small eigenvalues. Mathematically, by adding a diagonal matrix ,which is strictly non-singular to the original $A^{T}A$ as shown in Equ.19, we obtain matrix $\tilde{A}$ with larger eigenvalues. Intuitively, one extreme way to decrease the value of $\beta$ would be to make splats either transparent, or fully opaque. In this extreme case, each row in $\tilde{A}$ would contain only a single non-zero entry with a value of one, yielding a globally optimal solution. We employed such regularization by carefully adjusting the opacity activation term during feature lifting. Compare to a linear adjustment like original Tikhonov Regularizer, our method utilizes a non-linear soft guidance without undermine the visual quality. We introduce more implementation details about the Tikhonov Guidance in the appendix.A.
@@
Table 1: IoU comparison on LeRF OVS. The symbol $\dagger$ denotes results reported directly from the original papers, while $\ddagger$ indicates results evaluated using their official implementations. Our mIoU evaluation follows the LAGA(Cen et al. 2025) implementation. For a fair comparison, we use the same 3D Gaussian model as the lifting primitive, trained using the official repository provided by LAGA(Cen et al. 2025). The same trained model is used for LAGA, DrSplat, Occam’LGS, and our method in all $\ddagger$ evaluations. F denotes Figurines, T is Teatime, R is Ramen, W is Waldo kitchen. We use SAM+OpenCLIP for encoding for fair comparison
Method/Scene F. T. R. W. Mean
$\text{LSeg}^{\dagger}$ (Li et al. 2022) 7.6 21.7 7.0 29.9 16.6
$\text{LeRF}^{\dagger}$ (Kerr et al. 2023) 38.6 45.0 28.2 37.9 37.4
$\text{N2F2}^{\dagger}$ (Bhalgat et al. 2024) 47.0 69.2 56.6 47.9 54.4
$\text{LangSplat}^{\dagger}$ (Qin et al. 2024) 25.9 35.6 29.3 33.5 31.1
$\text{LeGaussian}^{\dagger}$  (Shi et al. 2024) 31.2 34.5 17.6 17.3 25.2
$\text{SuperGseg}^{\dagger}$ (Liang et al. 2024) 43.7 55.3 18.1 26.7 35.9
$\text{VLGS}^{\dagger}$ (Peng et al. 2024a) 58.1 73.5 61.4 54.8 62.0
$\text{SAGA}^{\dagger}$ (Cen et al. 2023) 36.2 19.3 53.1 14.4 30.7
$\text{OpenGaussian}^{\dagger}$ (Wu et al. 2025) 61.1 59.1 29.2 31.9 45.3
$\text{GS Grouping}^{\dagger}$ (Ye et al. 2023) 60.9 40.0 45.5 38.7 46.3
$\text{LAGA}^{\ddagger}$ (Cen et al. 2025) 56.1 68.9 57.4 64.6 61.7
$\text{LAGA}^{\dagger}$ (Cen et al. 2025) 64.1 70.9 55.6 65.6 64.0
$\text{DrSplat}^{\ddagger}$ (Jun-Seong et al. 2025) 47.5 66.2 36.7 47.5 49.5
$\text{DrSplat}^{\dagger}$ (Jun-Seong et al. 2025) 53.4 57.2 24.7 39.1 43.6
$\text{OccamLGS}^{\ddagger}$ (Cheng et al. 2024) 60.1 68.3 55.3 47.7 57.8
$\text{OccamLGS}^{\dagger}$ (Cheng et al. 2024) 58.6 70.2 51.0 65.3 61.3
Ours 67.6 68.5 62.3 62.1 65.1
Post Aggregation Filtering
Inspired by LAGA (Cen et al. 2025), we cluster the lifted features to identify and remove noise from SAM‐generated masks, as shown as Equ.20a, where -1 denotes unclassified splats, and $K+1$ is the number of clusters. Unlike LAGA, which learns separate affinity features, we simply reuse the Tikhonov‐Guided solution $\tilde{x}$ from Eq. 19 as our clustering feature. This immediately assigns each splat to a cluster. We then encode each Gaussian’s cluster ID as a one‐hot signal and render this signal back to 2D. By applying an argmax operation, we recover a 2D mask for each cluster, establishing a pixel‐level correspondence between the clusters and the original SAM masks as shown in Equ.20b and Equ.20c (Note: Any orthogonal encoding-decoding System could be used here).
The remaining masks—denoted as $B^{\prime}$—are significantly more self‐consistent and better aligned with the underlying decomposition of splat primitives. The visualization of the alignment is displayed in the technical appendix.B. This produces a well–posed data set that further boosts downstream semantic‐segmentation and object‐querying performance to state‐of‐the‐art levels. In contrast to LAGA’s complex view‐dependent clustering, our simpler Post Aggregation Filtering approach reveals that most apparent “view‐dependent” variations stem from mask noise rather than useful multi‐view information. Noisy masks can be found in Fig.1 as well as Fig.9. We further provide many mis-leading masks’ visualization and trust worthy masks in the appendix.B.
$\gamma$ $=\mathrm{Agg}\bigl{(}x\bigr{)}\;\in\{-1,\dots,K\}^{P},$ (20a)
$\Gamma$ $=\mathrm{onehot}(\gamma)\;\in\{0,1\}^{P\times(K+1)},$ (20b)
$\kappa$ $=\arg\max\bigl{(}A\,\Gamma\bigr{)}-1\;\in\{-1,\dots,K\}^{R},$ (20c)
@@
$M(j,B)$ $=m_{j}\in\mathbb{R}^{H\times W},M(j,\kappa)=m^{\prime}_{j}\in\mathbb{R}^{H\times W}$ (21)
$B^{\prime}_{j}$ $=\begin{cases}B_{j},&\text{IoU }\bigl{(}m_{j},m^{\prime}_{j}\bigr{)}\geq\tau,\\[6.0pt] \emptyset,&\text{otherwise.}\end{cases}$ (22)
Auto Threshold Selection
@@
We observe that threshold selection can significantly alter the final result. Rather than manually choosing thresholds per object (as in (Qin et al. 2024)), we leverage the high contrast of the raw attention map for automatic selection. As shown in Fig. 11, even though the ideal hard threshold for correctly isolating chopsticks shifts from 0.225 to 0.30 in the raw attention values, the first valley in the attention histogram remains clearly identifiable. We therefore derive our threshold directly from the histogram’s local extrema. We first locate its largest peak and then take the adjacent, smaller valley as the threshold. Further details are provided in the Appendix.H
Experiment
Figure 2: Qualitative comparison of semantic masks between our method compare to DrSplats on ramen scene in LERF-OVS. Our method produces more accurate masks aligning with the objects in the scene. More qualitative result could be found in appendix.E
Although our solver can lift any 2D feature into high-dimensional primitives, few benchmarks exist to evaluate such capability. We evaluate our approach on the LeRF-OVS (Qin et al. 2024) and 3D-OVS (Liu et al. 2023) datasets, using the segmentation masks annotated from (Qin et al. 2024; Kerr et al. 2023), and report mean Intersection-over-Union (mIoU) according to the protocols of (Cen et al. 2025; Jun-Seong et al. 2025). As Dr. Splat’s official code was not available at the time of writing, we follow the evaluation protocol provided by LAGA for consistency and comparability. CosegGS and Occam’s LGS (Dou et al. 2024; Cheng et al. 2024), both based on the same naïve solver (Eq. 7), are represented here by a single baseline. As shown in Tab.1, we group the methods into four categories: 2D- and NeRF-based, training-based, grouping-based, and heuristic forward-lifting. First, second, and third place in the benchmark are highlighted in light red, light pink, and light yellow, respectively. Our results for the 3D-OVS(Liu et al. 2023) dataset is provided in the appendix.E
Table 2: Multi Kernel Comparison on mIoU. For this comparison, we use our method without Tikhonov-Guidance, Post Aggregation Filtering, or auto threshold selection to illustrate its kernel agnostic capability. We evaluate performance using DBS, 2DGS, as well as 3DGS, # represents the inria(Kerbl et al. 2023) implementation, * represents Gsplat(Tancik et al. 2023) implementation
@@
DBS (Liu et al. 2025) 49.5 50.8 51.2 61.3 53.2
$\text{3DGS}^{*}$ (Tancik et al. 2023) 55.3 63.5 47.8 49.8 54.1
$\text{3DGS}^{\#}$ (Kerbl et al. 2023) 60.1 68.3 55.3 47.7 57.8
$\text{2DGS}^{*}$ (Huang et al. 2024) 62.0 66.3 56.0 51.1 58.9
We demonstrate our kernel-agnostic capability with a multi-kernel comparison Tab.2. In this study, the 2DGS implementation outperforms all other representations due to its specially designed kernel that encourages high opacity values, yielding a naturally lower $\beta$ as a result. For meaningful PSNR computation, we scale each final rendered feature to the range [1,1][-1,1] and normalize it to unit Euclidean norm. Consequently, the peak signal in the MSE loss is 4, the square of the maximum possible error (i.e., 2).
To compare feature lifting quality more broadly, we project a variety of descriptors(Fu et al. 2024)—MaskCLIP, CLIP, ViT, DINO, and ResNet—through our solver (prior to any post-aggregation filtering) and visualize the results using cosine similarity and PCA (see the appendix.G for details). We avoid PSNR for evaluation since MSE-based metrics scale with the number of feature channels. For example, DINO features (384 channels) yield consistently lower PSNR than that of CLIP features (512 channels) despite having comparable perceptual quality. Instead, we demonstrate both quantitatively via cosine similarity as mentioned in Tab.3, and qualitatively via PCA visualization that our lifted projections consistently outperform the raw descriptors, making direct PSNR comparisons potentially misleading. While we conduct experiments with different types of primitives—including 2DGS, 3DGS, and DBS, for comparing feature descriptors we use only 3DGs as the standard primitive to ensure fairness.
Table 3: Multi Feature Comparison on Cosine-Similarity. We evaluate the feature-agnostic capability without Tikhonov guidance, post-aggregation filtering, or automatic threshold selection. We apply our method to dense features generated by FeatUp(Fu et al. 2024), including features from DINO, DINOv2, CLIP, MaskCLIP, ViT, and ResNet, using the Inria implementation. For a fair comparison, we use geometry provided by Gsplat.
Features/scene F. T R W. Mean
SAMOpenCLIP (Kirillov et al. 2023) 89.3 90.7 90.6 91.0 90.4
MaskCLIP (Dong et al. 2023) 92.4 94.0 94.5 94.7 93.9
CLIP (Radford et al. 2021) 91.9 93.7 94.6 94.9 93.8
DINO (Caron et al. 2021) 78.7 80.2 83.0 82.0 81.0
DINOv2 (Oquab et al. 2023) 83.5 85.6 89.6 89.2 87.0
ViT (Dosovitskiy et al. 2020) 84.6 83.7 86.4 88.0 85.7
ResNet (He et al. 2016) 95.8 94.8 96.2 97.0 96.0
Table 4: Runtime Comparison. We compare the runtime of our lifting method with LAGA and DrSplats, along with the ability to preserve the full feature dimensionality. Here $T_{i}$ denotes Tikhonov Guidance, P denotes Post Aggregation Filtering. All timings are measured on a single RTX 4090 GPU using CUDA Toolkit 12.8 for consistency. Notice that the runtime of lifting-based methods scales linearly with the number of views, whereas training-based and grouping-based methods are determined by predefined training steps. For DrSplats, as its computation time scales with both the number of views and the Top-K value, we report results using Top-40 only.
Features/scene F. T R W. Mean
@@
To compare runtimes, we report wall-clock times for lifting, and post-processing using standard SAM+OpenCLIP as the encoder as mentioned in Tab.4. Additionally, we perform further experiments using MaskCLIP as an encoder, which are included in the appendix.D. The motivation behind this choice is that standard SAM+OpenCLIP typically takes approximately one hour to pre-process the dataset, whereas MaskCLIP completes the same pre-processing task within minutes. Additional mIoU results for MaskCLIP are also provided in the appendix.D. The splats training time per scene is roughly 20 minutes for the Gsplats (Tancik et al. 2023) training pipeline and approximately 40 minutes for the Inria-based pipeline. Regarding memory consumption, our implementation can lift features with over 512 channels without encountering CUDA out-of-memory (OOM) errors. In contrast, in our testing environment, DrSplat’s implementation could only handle up to 32 channels before hitting memory limitations. Lifting-based methods generally exhibit linear complexity with respect to the number of feature channels, the number of view images, and the number of witness splats per ray.
Finally, our ablation study isolates the impact of each component—naïve preconditioner, full Tikhonov (as shown in Table 5)—and post-aggregation filtering. We also examine implementation details, including feature scaling, multi-level SAM mask integration, and segmentation-threshold selection. All qualitative results and additional engineering notes are provided in the appendix.D
Table 5: Ablation Study. We evaluate the impact of each module on the LeRF data set in a step-by-step manner. $T_{i}$ denotes Tikhonov Guidance, P denotes Post Aggregation Filtering, and A represents Automatic threshold selection. The baseline corresponds to the row-sum formulation in Equ.24
Features/scene F. T R W. Mean
Ours w/o ($T_{i}$PA) 60.1 68.3 55.3 47.7 57.8
Ours w/o (PA) 61.7 67.8 53.8 49.6 58.2
Ours w/o ($T_{i}$A) 65.5 72.0 58.6 50.4 61.6
Ours w/o (A) 64.8 71.6 61.7 54.7 63.2
Ours 67.6 62.3 68.5 62.1 65.1
Discussion

### v2

@@ 3.2 Solver
Here, we start with the original loss function $\mathcal{L}(x)$, and calculate a surrogate loss function $\mathcal{J}(x)$. The norm $||\cdot||$ represents any convex loss function such as L1, L2, or a Huber loss. From Property.2, we apply the Jensen’s inequality. This means the surrogate loss function $\mathcal{J}$ is larger than $\mathcal{L}$ and serves as an upper bound for the true loss. For the sake of argument, we consider a special case where $||\cdot||$ is the L2 loss. If $\mathcal{J}$ attains a minimum, it occurs where the gradient is zero, as shown in Equ.9. Therefore, we obtain an optimal solution on the surrogate loss function $\mathcal{J}$.
$\frac{\partial\mathcal{J}}{\partial x_{j}}=\sum_{i}A_{ij}\left(x_{j}-B_{i}\right),\quad\frac{\partial\mathcal{J}}{\partial x_{j}}=0\quad\Rightarrow\quad x^{\prime}_{j}=\frac{\sum_{i}A_{ij}B_{i}}{\sum_{i}A_{ij}}$ (9)
We now define $\beta$, which intuitively measures the dispersion of the lifted feature along a viewing ray at the global optimal lift. Suppose the optimal lift yields a solution $\hat{x}$. We define $\beta$ at Equ.11.
@@
Notice that usually $x^{\prime}\neq\hat{x}$. Therefore, $x^{\prime}$ is not optimal in $\mathcal{J}$. By Equ.8, we have the following equation under the L2 loss:
@@
From the analysis, we can conclude that the methods inDou et al. (2024); Cheng et al. (2024) are also bounded by the $1+\beta$ term. The method Chacko et al. (2025) is bounded when a $l_{\infty}$ norm function is applied. The approach in Joseph et al. (2024) is bounded only when the input observations are unit vectors. Jun-Seong et al. (2025) does not provide a theoretical bound, as it selects the top-K splats based solely on depth ordering.
Figure 2: Qualitative comparison on the LeRF-OVS Ramen scene. We compare our method against DrSplats and the Ground Truth. As shown in the legend, distinct colors represent different semantic classes (e.g., egg, chopsticks). Our method performs better compared to recent SOTA Jun-Seong et al. (2025). More qualitative result could be found in Fig. 3, Fig. 4, and Fig. 5
Although our solver can lift any 2D feature into high-dimensional primitives, few benchmarks exist to evaluate such capability. We evaluate our approach on the LeRF-OVS Qin et al. (2024) and 3D-OVS Liu et al. (2023) datasets, using the segmentation masks annotated from Qin et al. (2024); Kerr et al. (2023), and report mean Intersection-over-Union (mIoU) according to the protocols of Cen et al. (2025); Jun-Seong et al. (2025). As Dr. Splat’s official code was not available at the time of writing, we follow the evaluation protocol provided by LAGA for consistency. CosegGS and Occam’s LGS Dou et al. (2024); Cheng et al. (2024), both based on the same naïve solver (Eq. 6), are represented here by a single baseline. As shown in Tab.1, we group the methods into four categories: 2D- and NeRF-based, training-based, grouping-based, and heuristic forward-lifting. First, second, and third place in the benchmark are highlighted in light red, light pink, and light yellow, respectively. Results for the 3D-OVSLiu et al. (2023) dataset is provided in the appendix.F
3.3 Regularizer
@@
3.3.1 Tikhonov Guidance
According to property.4 and inspired by Tikhonov and Arsenin (1977); Tikhonov (1963), we observe that emphasizing diagonal dominance helps mitigate noise from linear system A. Therefore, we propose a carefully designed regularizer applied during solving without sacrificing the RGB rendered result. Specifically, we first convert Equ.6 into a fully diagonal row-sum version. . Briefly speaking, stabilizing the linear system $A^{T}A$ involves avoiding small eigenvalues. Mathematically, by adding a diagonal matrix ,which is strictly non-singular to the original $A^{T}A$ as shown in Equ.14, we obtain matrix $\tilde{A}$ with larger eigenvalues. Intuitively, one extreme way to decrease the value of $\beta$ would be to make splats either transparent, or fully opaque. In this extreme case, each row in $\tilde{A}$ would contain only a single non-zero entry with a value of one, yielding a globally optimal solution. We employed such regularization by carefully adjusting the opacity activation term during feature lifting. Compare to a linear adjustment like original Tikhonov Regularizer, our method utilizes a non-linear soft guidance without undermine the visual quality. We introduce more implementation details about the Tikhonov Guidance in the appendix.B.
@@ 3.3.1 Tikhonov Guidance
3.3.2 Post Aggregation Filtering
Inspired by LAGA Cen et al. (2025), we cluster the lifted features to identify and remove noise from SAM‐generated masks, as shown as Equ.15, where -1 denotes unclassified splats, and $K+1$ is the number of clusters. Unlike LAGA, which learns separate affinity features, we simply reuse the Tikhonov‐Guided solution $\tilde{x}$ from Eq. 14 as our clustering feature. This immediately assigns each splat to a cluster. We then encode each Gaussian’s cluster ID as a one‐hot signal and render this signal back to 2D. By applying an argmax operation, we recover a 2D mask for each cluster, establishing a pixel‐level correspondence between the clusters and the original SAM masks as shown in Equ.15 and Equ.15 (Note: Any orthogonal encoding-decoding System could be used here).
The remaining masks—denoted as $B^{\prime}$—are significantly more self‐consistent and better aligned with the underlying decomposition of splat primitives. The visualization of the alignment is displayed in the technical appendix.C. This produces a well–posed data set that further boosts downstream semantic‐segmentation and object‐querying performance to state‐of‐the‐art levels. In contrast to LAGA’s complex view‐dependent clustering, our simpler Post Aggregation Filtering approach reveals that most apparent “view‐dependent” variations stem from mask noise rather than useful multi‐view information. Noisy masks can be found in Fig.1 as well as Fig.7. We further provide many mis-leading masks’ visualization and trust worthy masks in the appendix.C.
$\gamma=\mathrm{Agg}\bigl(x\bigr)\;$ $\in\{-1,\dots,K\}^{P},\quad\Gamma=\mathrm{onehot}(\gamma)\;\in\{0,1\}^{P\times(K+1)},$ (15a)
$\kappa=\arg\max\bigl(A\,\Gamma\bigr)-1\;\in\{-1,\dots,K\}^{R},$ (15b)
@@ 3.3.2 Post Aggregation Filtering
[…]
3.3.3 Auto Threshold Selection
@@
We observe that threshold selection can significantly alter the final result. Rather than manually choosing thresholds per object (as in Qin et al. (2024)), we leverage the high contrast of the raw attention map for automatic selection. As shown in Fig. 15, even though the ideal hard threshold for correctly isolating chopsticks shifts from 0.225 to 0.30 in the raw attention values, the first valley in the attention histogram remains clearly identifiable. We therefore derive our threshold directly from the histogram’s local extrema. We first locate its largest peak and then take the adjacent, smaller valley as the threshold. Further details are provided in the Appendix.I
4 Experiment
Figure 3: Qualitative Comparison of Attention Maps and Segmentation Masks. We present a qualitative comparison between our proposed method (top row) and the baseline, Dr. Splat (middle row). The heatmaps visualize the model’s attention response, or affinity feature projection, to various text prompts (e.g., ”Green apple,” ”Jake the dog,” ”Miffy”). The bottom row compares the final semantic segmentation masks for the entire scene against the ground truth.
Table 1: LeRF OVS (left) vs. Multi-Feature & Ablation (right).
(a)IoU comparison on LeRF OVS. The symbol $\dagger$ denotes results reported directly from the original papers, while $\ddagger$ indicates results evaluated using their official implementations. Our mIoU evaluation follows the LAGACen et al. (2025) implementation. For a fair comparison, we use the same 3D Gaussian model as the lifting primitive, trained using the official repository provided by LAGACen et al. (2025). The same trained model is used for LAGA, DrSplat, Occam’LGS, and our method in all $\ddagger$ evaluations. F denotes Figurines, T is Teatime, R is Ramen, W is Waldo kitchen. We use SAM+OpenCLIP for encoding for fair comparison
Method/Scene F T. R. W. Mean
$\text{LSeg}^{\dagger}$ Li et al. (2022) 7.6 21.7 7.0 29.9 16.6
$\text{LeRF}^{\dagger}$ Kerr et al. (2023) 38.6 45.0 28.2 37.9 37.4
[…]
$\text{LangSplat}^{\dagger}$ Qin et al. (2024) 25.9 35.6 29.3 33.5 31.1
$\text{LeGaussian}^{\dagger}$  Shi et al. (2024) 31.2 34.5 17.6 17.3 25.2
$\text{SuperGseg}^{\dagger}$ Liang et al. (2024) 43.7 55.3 18.1 26.7 35.9
[…]
$\text{SAGA}^{\dagger}$ Cen et al. (2023) 36.2 19.3 53.1 14.4 30.7
[…]
$\text{GS Grouping}^{\dagger}$ Ye et al. (2023) 60.9 40.0 45.5 38.7 46.3
[…]
$\text{DrSplat}^{\ddagger}$ Jun-Seong et al. (2025) 47.5 66.2 36.7 47.5 49.5
$\text{DrSplat}^{\dagger}$ Jun-Seong et al. (2025) 53.4 57.2 24.7 39.1 43.6
$\text{OccamLGS}^{\ddagger}$ Cheng et al. (2024) 60.1 68.3 55.3 47.7 57.8
[…]
(b)Multi Feature Comparison on Cosine-Similarity. We evaluate the feature-agnostic capability without Tikhonov guidance, post-aggregation filtering, or automatic threshold selection. We apply our method to dense features generated by FeatUp Fu et al. (2024), including features from DINO, DINOv2, CLIP, MaskCLIP, ViT, and ResNet. For a fair comparison, we use geometry provided by Gsplat to obtain the geometry for all scenes. Notice that SAM + OpenCLIP is follow exactly the LangSplatQin et al. (2024) implementation.
Features/Scene F. T R W. Mean
SAM+OpenCLIP Kirillov et al. (2023) 89.3 90.7 90.6 91.0 90.4
MaskCLIP Dong et al. (2023) 92.4 94.0 94.5 94.7 93.9
CLIP Radford et al. (2021) 91.9 93.7 94.6 94.9 93.8
DINO Caron et al. (2021) 78.7 80.2 83.0 82.0 81.0
DINOv2 Oquab et al. (2023) 83.5 85.6 89.6 89.2 87.0
ViT Dosovitskiy et al. (2020) 84.6 83.7 86.4 88.0 85.7
ResNet He et al. (2016) 95.8 94.8 96.2 97.0 96.0
(c)Ablation Study. We evaluate the impact of each module on the LeRF data set in a step-by-step manner. $T_{i}$ denotes Tikhonov Guidance, P denotes Post Aggregation Filtering, and A represents Automatic threshold selection. The baseline corresponds to the row-sum formulation in Equ.18. To further understand which components is effective We provide a full ablation study shown in the appendix, Tab.4.
@@ 4 Experiment
Ours w/o ($T_{i}$PA) 60.1 68.3 55.3 47.7 57.8
Ours w/o (PA) 61.7 67.8 53.8 49.6 58.2
Ours w/o ($T_{i}$A) 65.5 72.0 58.6 50.4 61.6
Ours w/o (A) 64.8 71.6 61.7 54.7 63.2
Ours (full) 67.6 62.3 68.5 62.1 65.1
Table 2: Multi-Kernel and Runtime Comparison. Left: Multi-Kernel across different splats. Right: end-to-end runtime.
(a)Multi Kernel Comparison on mIoU. For this comparison, we use our method without Tikhonov-Guidance, Post Aggregation Filtering, or auto threshold selection to illustrate its kernel agnostic capability. We evaluate performance using DBS, 2DGS, as well as 3DGS, # represents the inriaKerbl et al. (2023) implementation tuned hyper parameters exactly according to LAGACen et al. (2025) released code, * represents GsplatTancik et al. (2023) implementation with no additional hyper parameters changed except set the background color to random.
Method F. T R W. Mean
DBS Liu et al. (2025) 49.5 50.8 51.2 61.3 53.2
$\text{3DGS}^{*}$ Tancik et al. (2023) 55.3 63.5 47.8 49.8 54.1
$\text{3DGS}^{\#}$ Kerbl et al. (2023) 60.1 68.3 55.3 47.7 57.8
$\text{2DGS}^{*}$ Huang et al. (2024) 62.0 66.3 56.0 51.1 58.9
(b)Runtime Comparison. We compare the runtime of our lifting method with LAGA and DrSplats, along with the ability to preserve the full feature dimensionality. Here $T_{i}$ denotes Tikhonov Guidance, P denotes Post Aggregation Filtering. All timings are measured on a single RTX 4090 GPU using CUDA Toolkit 12.8 for consistency. Notice that the runtime of lifting-based methods scales linearly with the number of views, whereas training-based and grouping-based methods are determined by predefined training steps. For DrSplats, we report results using Top-40 follow their implementation.
Method F. T R W. Mean
@@
We demonstrate our kernel-agnostic capability with a multi-kernel comparison Tab.2. In this study, the 2DGS implementation outperforms all other representations due to its specially designed kernel that encourages high opacity values, yielding a naturally lower $\beta$ as a result.
For downstream segmentation tasks, qualitative results in Fig.2, Fig.5, Fig. 3 and Fig. 4 demonstrate that our method generates significantly clearer and more localized attention maps for object queries. In contrast, while the baseline (Dr. Splat) identifies the general location of objects, its activations are more diffuse and less uniform. This lack of precision leads to lower performance in downstream segmentation. As observed, our focused attention maps translate into segmentation masks that closely match the ground truth, exhibiting fewer false positives and less noise than the baseline. These qualitative observations align well with the quantitative mIoU results reported in Table 1.
To compare feature lifting quality more broadly, we project a variety of descriptorsFu et al. (2024)—MaskCLIP, CLIP, ViT, DINO, and ResNet—through our solver (prior to any post-aggregation filtering) and visualize the results using cosine similarity and PCA. Tab.9, Fig.13 and Fig.14 show consistent improvement on current SOTA lifting methodsGuo et al. (2024); Chacko et al. (2025). (see the appendix.H for details). We avoid PSNR for evaluation since MSE-based metrics scale with the number of feature channels. For example, DINO features (384 channels) yield consistently lower PSNR than that of CLIP features (512 channels) despite having comparable perceptual quality. Instead, we demonstrate both quantitatively via cosine similarity as mentioned in Tab.1, and qualitatively via PCA visualization that our lifted projections consistently outperform the raw descriptors, making direct PSNR comparisons potentially misleading. While we conduct experiments with different types of primitives—including 2DGS, 3DGS, and DBS, for comparing feature descriptors we use only 3DGs as the standard primitive to ensure fairness.
To compare runtimes, we report times for lifting, and post-processing using standard SAM+OpenCLIP as the encoder as mentioned in Tab.2. Additionally, we perform further experiments using MaskCLIP as an encoder, which are included in the appendix.E. The motivation behind this choice is that standard SAM+OpenCLIP typically takes approximately one hour to pre-process the dataset, whereas MaskCLIP completes the same pre-processing task within minutes. Additional mIoU results for MaskCLIP are also provided in the appendix.E. The splats training time per scene is roughly 20 minutes for the Gsplats Tancik et al. (2023) training pipeline and approximately 40 minutes for the Inria-based pipeline. Regarding memory consumption, our implementation can lift features with over 512 channels without encountering CUDA out-of-memory (OOM) errors. In contrast, in our testing environment, DrSplat’s implementation could only handle up to 32 channels before hitting memory limitations. Lifting-based methods exhibit linear complexity with respect to the number of feature channels. , view images, and witness splats per ray.
Finally, our ablation study isolates the impact of each component—naïve preconditioner, full Tikhonov (as shown in Table 4)—and post-aggregation filtering. We also examine implementation details, including feature scaling, multi-level SAM mask integration, and segmentation-threshold selection. All qualitative results and additional engineering notes are provided in the appendix.E
5 Discussion